A small plane leaves an airport and flies due south at 240km/h. A jet leaves the same airport 2 hours later and flies due south at 960km/h. How long does it take the jet to catch up to get to the small plane
CAN YOU SHOW HOW YOU GOT IT..

Answer by Alan3354(69443)

A small plane leaves an airport and flies due south at 240km/h. A jet leaves the same airport 2 hours later and flies due south at 960km/h. How long does it take the jet to catch up to get to the small plane
------------
In 2 hours, the 1st plane is 480 km away (240*2)
The 2nd plane gains on the first at 720 km/hr (960 - 240)
-----
d = r*t
t = d/r
480/720 = 2/3 hours to catch up